Circular No. 5828

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 MF The following improved orbital elements are from 34 observations, June

22-30. The additional observations will be published on the Aug. 2 MPCs.

T = 1993 Aug. 1.322 TT  ω = 74.^{o}639
e = 0.53971  Ω = 241.316 }2000.0
q = 1.14629 AU  i = 8.114
a = 2.49035 AU n^{o} = 0.250791 P = 3.93 years
1993TT α_{2000} δ_{2000} Δ r ϵ β V
July 2 20^{h}43.^{m}02 +12^{o}27.^{′}8 0.245 1.200 134.^{o}1 37.^{o}4 13.4
4 20 49.75 +14 29.6 0.240 1.193 132.9 38.6 13.3
6 20 56.64 +16 31.2 0.237 1.187 131.6 39.8 13.3
8 21 03.69 +18 31.8 0.234 1.181 130.3 41.0 13.3
10 21 10.89 +20 30.6 0.231 1.176 129.1 42.2 13.3
12 21 18.22 +22 26.7 0.229 1.171 127.8 43.3 13.3
14 21 25.67 +24 19.3 0.228 1.166 126.5 44.4 13.3
16 21 33.22 +26 07.8 0.227 1.162 125.3 45.5 13.3
18 21 40.87 +27 51.4 0.226 1.159 124.2 46.5 13.4
20 21 48.58 +29 29.7 0.226 1.155 123.2 47.4 13.4
22 21 56.33 +31 02.1 0.226 1.153 122.2 48.2 13.4
24 22 04.10 +32 28.3 0.227 1.150 121.4 48.9 13.4
26 22 11.88 +33 48.1 0.228 1.149 120.6 49.5 13.4

SUPERNOVA 1993J IN NGC 3031 S. D. Van Dyk, H.-A. Nguyen, and K. W. Weiler, Naval Research

Laboratory; R. A. Sramek and M. P. Rupen, National Radio Astronomy Observatory; and N. Panagia, Space Telescope Science Institute, report that new observations of SN 1993J made with the Very Large Array in ‘CnB’ and ‘C’ configurations show that radio emission from the supernova is now detectable at 20-cm wavelength (1.45 GHz) for the first time. On and prior to 1993 May 25, no emission was detectable to a limit (3σ) of 0.25 mJy in that band, while by June 11, 18, and 25, the source was detectable at levels of 1.3, 1.8, and 3.6 mJy, respectively. The source remains strong (> 50 mJy) in the three previously detected radio bands (1.3, 2, and 3.6 cm) and continues to rapidly increase in flux density at 6 cm. Monitoring with the VLA in all bands continues.

1993 July 2 Daniel W. E. Green